OpenStudy (anonymous):

Am I wrong Find the equation of the line passing through the points (-6, 6) and (-3, -4) My Answer: y=30+4x

OpenStudy (loser66):

show your work,please

OpenStudy (anonymous):

(-4) - 6 / (-3) - (-6) = -12/3 = -4

OpenStudy (anonymous):

hold up my calculation was wrong its -10/3

OpenStudy (loser66):

ok, you are good, continue

OpenStudy (anonymous):

y1-y= -10/3 (x1 - X)

OpenStudy (loser66):

stop

OpenStudy (anonymous):

?

OpenStudy (loser66):

you have m, you have a point, let pick (-6,6) just put them on y = mx +b to find b

OpenStudy (anonymous):

so my slope is -10/3?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

so y=-10/3+b

OpenStudy (loser66):

ok, for easier, let pick (3, -4)

OpenStudy (loser66):